# Benchmarking Performance, Explainability, and Evaluation Strategies of Vision-Language Models for Surgery: Challenges and Opportunities

### v1

@@ 3.2.1 Grad-CAM for ResNest VLM
For VLM that use residual net Yuan et al. (2023, 2024b, 2024a), the Grad-Cam formula is defined as following. Let $A\in\mathbb{R}^{K\times H\times W}$ denote the output of a selected convolutional layer, where $K$ is the number of channels, and $H\times W$ is the spatial resolution. Let $A_{k}(i,j)$ denote the activation at position $(i,j)$ in channel $k$. The target output score is denoted by $S_{c}\in\mathbb{R}$, which in the case of VLMs is typically a similarity score between image and text. Let $S_{c}\in\mathbb{R}$ denote the output score corresponding to the model’s prediction or the score of interest. In the case of vision-language models, $S_{c}$ is often defined as the cosine similarity between the image embedding $\phi_{\text{img}}(I)$ and the text embedding $\phi_{\text{text}}(T_{c})$ of a selected class $c$:
$S_{c}=\frac{\phi_{\text{img}}(I)^{\top}\phi_{\text{text}}(T_{c})}{\|\phi_{\text{img}}(I)\|\cdot\|\phi_{\text{text}}(T_{c})\|}.$
Finally, computing similarity scores for a set of candidate text prompts $\{T_{1},T_{2},\dots,T_{N}\}$ and select the one with the highest score:
$\hat{c}=\max_{c\in\{1,\dots,N\}}S_{c},$
where $\hat{c}$ is the predicted class, and Let $S_{\hat{c}}$ denotes the scalar target used for backpropagation in CAM methods.
Grad-CAM computes the gradient of the score $S_{c}$ with respect to the feature map $A$, applies global average pooling to obtain channel importance weights:
$\alpha_{k}=\frac{1}{H\cdot W}\sum_{i=1}^{H}\sum_{j=1}^{W}\frac{\partial S_{\hat{c}}}{\partial A_{k}(i,j)},$
The gradient tells how much each feature map channel $A_{k}$ contributes to the model’s prediction $S_{\hat{c}}$. it does this by averaging the gradient of predicted score with respect to all locations $(i,j)$ in that channel and use it act as a weight that reflects the importance of that channel. Based on it, we can then construct the heatmap as:
$L_{\text{Grad-CAM}}(i,j)=\text{ReLU}\left(\sum_{k=1}^{K}\alpha_{k}\cdot A_{k}(i,j)\right).$ (1)
For VLM that use transformer, we sue the following method.
@@ 3.2.2 Gradient-Based Attention Rollout for CLIP-ViT
Grad-CAM can be adapted to Vision Transformers used in vision-language models (VLMs), specifically CLIP Radford et al. (2021) with a ViT image encoder. Following the method from Chefer et al. (2021), instead of using convolutional feature maps, we leverage the self-attention matrices and their gradients to identify input image patches that contribute most to the similarity between an image and a given text prompt.
Let $\mathcal{T}^{(l)}\in\mathbb{R}^{N\times N}$ denote the self-attention matrix from the $l$-th transformer block in the ViT encoder, where $N$ is the number of tokens, including the [CLS] token. We extract $\mathcal{T}^{(l)}$ after the softmax operation during the forward pass.
@@
$\nabla\mathcal{T}^{(l)}=\frac{\partial S_{\hat{c}}}{\partial\mathcal{T}^{(l)}},$
where $S_{\hat{c}}$ is the predicted similarity score corresponding to the selected text prompt in a zero-shot setting. To capture influence on the similarity score, we compute the gradient-weighted attention map:
@@
We reshape and upsample the vector $[r_{1},\dots,r_{N-1}]\in\mathbb{R}^{N-1}$ into the input image resolution using bilinear interpolation. The final heatmap is defined as:
$L_{\text{GradCAM-ViT}}(i,j)=\text{Upsample}\left(\text{reshape}\left([r_{1},\dots,r_{N-1}]\right)\right)_{i,j},$ (2)
where $(i,j)$ indexes pixel locations in the original image. Based on the CAM style methods we now formally define our metrics as following.
3.3 Instrument classification metrics
Notation.
Let $I$ denote an input image of size $H\times W$, and let $T$ be a corresponding text prompt. The model produces a Grad-CAM heatmap $L(i,j)\in\mathbb{R}^{H\times W}$ using (1) or (2), where each pixel $(i,j)$ reflects the spatial importance of that location for the model’s prediction. We normalize $L$ to the range $[0,1]$, and apply a fixed threshold $\tau\in[0,1]$ to etain information about the "strength" of different regions. We now define the predicted $\tau$-attention region as following:
@@ 3.3 Instrument classification metrics
Let $\mathcal{C}=\{c_{1},\dots,c_{K}\}$ be the set of ground-truth class labels (e.g., surgical instruments) present in the image. For each class $c_{k}\in\mathcal{C}$, let $\mathcal{G}_{c_{k}}^{(r)}\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ denote the set of pixel coordinates in the $r^{\text{th}}$ annotated region of class $c_{k}$, for $r=1,\dots,N_{k}$. Let $\hat{c}$ be the class predicted by the model. We define the set of all annotated pixels (regardless of class) as:
@@
This ensures that only alignment with correctly predicted classes contributes positively to the score. Beyond conventional classification tasks, understanding the surgical action depicted in an image is also critically important in surgical data.
3.4 Triplet evaluation metrics
@@
In the triplet classification task, the traditional and straightforward way of evaluating performance is to compute the F1 classification score for each class. However, in this paper, beyond assessing classification accuracy, we also challenge the spatial reasoning ability of vision-language models (VLMs). When it comes to localizing the bounding box of an action verb, the problem becomes more complex. As discussed in Twinanda et al. (2022), the localization of the action region is derived using class activation maps (CAMs) or attention mechanisms learned by the model. In particular, when an instrument is actively performing an action, the action is spatially grounded near the location of the instrument tip. Due to the lack of labeled action verb bounding boxes, we propose the Action Reasoning Score (ARS). This score measures whether the model’s attention for a predicted action (verb) overlaps with the bounding box of the predicted instrument (subject). Let a predicted triplet be denoted by $(s,v,o)$, where $s\in\mathcal{S}$ is the subject, $v\in\mathcal{V}$ is the verb (e.g., an action such as cut), and $o\in\mathcal{O}$ is the object (e.g., tissue) and ground-truth triplet is $(s^{*},v^{*},o^{*}).$We compute ARS only when: $s=s^{*}\quad\text{and}\quad v=v^{*}.$ Let the image be defined over a pixel grid of height $H$ and width $W$, with each pixel indexed by coordinates $(i,j)\in\{1,\dots,H\}\times\{1,\dots,W\}$. Let $L_{v}(i,j)\in[0,1]$ denote the value of Grad-CAM attention produced by (1) or (2) at pixel $(i,j)$ for the predicted verb $v$. We will later show how to generate the attention heatmap $L_{v}$ in section 4.1.
@@
Note that this score provides a soft measure of how much the verb-specific attention overlaps with the instrument’s spatial region, offering insight into whether the model is focusing on a plausible area for grounding the action. While there is no strict threshold for overlap, a higher value generally indicates better spatial alignment, whereas a near-zero value suggests the model is likely attending to an irrelevant region. A more detailed implementation is provided in the experimental analysis in Section 4.1.
(2) Triplet Classification:
[…]
Given a predicted triplet $(s,v,o)$ and a ground-truth triplet $(s^{*},v^{*},o^{*})$, we compute the proportion of correctly predicted components using:
In this work, we focus on evaluating the action reasoning ability of vision-language models (VLMs). Specifically, we are interested only in the classification of the subject and verb components, denoted by the indicators $\mathbbm{1}_{s=s^{*}}$ and $\mathbbm{1}_{v=v^{*}}$ in the ARS score (5).
@@ 4 Experiments Set up
4.1 Datasets and implementation
All experiments use a fixed Grad-CAM threshold of $\tau=0.3$ to generate binary attention maps for grounding evaluation. We first introduce all the datasets we used in this work:
Cholec80 BBox:
We use the Cholec80 BBox dataset Jalal et al. (2023); Alshirbaji et al. (2024), an extension of Cholec80 Twinanda et al. (2016b), which provides bounding box annotations for surgical instruments across 80 cholecystectomy videos. This dataset is used for the instrument classification task, allowing us to evaluate model predictions using spatially grounded metrics such as $\tau$-AC (3) and $\tau$-AA (4). It is also used as a source of spatial supervision when computing attention alignment for triplet classification in CholecT45.
CoPESD:
We also include the CoPESD dataset Wang et al. (2024), the first multimodal dataset for endoscopic submucosal dissection (ESD), which contains over 17,000 images with fine-grained, multi-level instrument annotations. This dataset is used for evaluating instrument classification and enables spatial reasoning analysis via the $\tau$-AC and $\tau$-AA metrics.
CholecT45:
[…]
@@ 4.1 Datasets and implementation
To align with the prompt-based input format of vision-language models (VLMs), we construct a pool of prompts using template-based formulations. For instrument classification, we use prompts template: “an image showing a {instrument class} in use”, covering all possible instrument classes in Cholec80. For triplet-based action reasoning, we use the template “I use a {instrument} to {verb} the {target}. In the instrument classification task, models are not required to predict all classes present in the image. Instead, we generate the heatmap $L_{\text{GradCAM}}$ and $L_{\text{GradCAM-ViT}}$ and evaluate spatial grounding using two metrics: $\tau$-AC, which assesses whether the model attends to any relevant region, and $\tau$-AA, which evaluates whether attention is correctly aligned with the specific object mentioned in the prompt.
For triplet reasoning, we compute the Action Reasoning Score (ARS) only when both the subject (instrument) and verb predictions are correct. In our implementation, similar to the instrument classification setup, we first generate a pool of triplet prompts using the template “I use a {instrument} to {verb} the {target}, covering all possible instrument, verb, and target classes. We then identify frames where both the predicted instrument and verb match the ground truth. For these correctly classified frames, we reuse the predicted verb to construct a secondary prompt of the form “I am performing {predicted_verb}. Finally, we generate its Grad-CAM heatmap $L_{v}$ and compute the action reasoning scores (5).
@@ 4.2 Benchmarked Models
In this work, we evaluate four vision-language models (VLMs): CLIP-ViTRadford et al. (2021), HecVLYuan et al. (2024b), PeskaVLPYuan et al. (2024a), and SurgVLPYuan et al. (2023). CLIP-ViT Radford et al. (2021) is the classic and widely adopted architecture for multi-modal learning, trained on large-scale natural image-text pairs. CLIP has demonstrated strong zero-shot performance across various domains, making it a natural reference point for evaluating domain transfer to surgical imagery.
The other three models—SurgVLPPeskaVLP, and HecVL are trained using surgical video datasets with varying levels of supervision, and architectural adaptations. These models are designed specifically to address the challenges of understanding visual-linguistic information in surgical environments, such as tool usage, procedural context, and anatomy. We aim to assess not only a deeper understanding of the reasoning capabilities of VLMs in surgical classification tasks, but also the limitations and trade-offs between general-purpose and task-specific VLMs within the current surgical domain.
Figure 2: Examples of triplet classification tasks using CLIP-ViT with Grad-CAM visualize the verb heatmap $L_{v}$, where CLIP-ViT successfully classifies all images in Panels (a) and (b) with the correct instrument and verb. Panels (a) illustrate Ideal action reasoning, with attention focused near the instrument and high ARS scores. In contrast, panels (b) show failure cases: despite correct triplet predictions, the attention is misplaced on background tissues, leading to ARS scores of 0.00. These examples highlight that some correct classifications are not based on the correct spatial evidences.
@@ 5.1 Instrument Classification Performance
Table 1 shows the spatial reasoning performance of five vision-language models across two datasets. Overall, all VLMs tested so far perform poorly under complex surgical environments. CLIP consistently outperforms the other models in both $\tau$-AC and $\tau$-AA metrics across most videos, achieving the highest $\tau$-AC of 0.3405 and $\tau$-AA of 0.0680 on the CoPESD dataset, and a $\tau$-AC of 0.2395 and $\tau$-AA of 0.1021 on the Cholec8k Bounding Box dataset. In contrast, despite being trained on surgical lectures, HecVL, PeskaVLP, and SurgVLP appear to focus more on understanding the overall surgical phase rather than performing fine-grained, frame-level instrument classification.

### v2

@@ 3.3.1 Grad-CAM for ResNest VLM
For VLMs that use ResNet-style backbones [9, 8, 7], the Grad-CAM formulation is defined as follows. Let $I\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 3}$ be the input image, and let $A\in\mathbb{R}^{K\times H\times W}$ denote the output of a selected convolutional layer in the image encoder, where $K$ is the number of channels and $H\times W$ is the spatial resolution. Let $A_{k}(i,j)$ denote the activation at spatial position $(i,j)$ in channel $k$. Let $\mathcal{P}=\{P_{1},P_{2},\dots,P_{N}\}$ be a set contain $N$ candidate natural language prompts. The vision-language model computes an embedding for the image $\phi_{\text{img}}(I)$ and for each text prompt $\phi_{\text{text}}(P_{k})$, and calculates a similarity score $S_{P_{k}}$ (typically cosine similarity) for each prompt in $\mathcal{P}$:
$S_{P_{k}}=\frac{\phi_{\text{img}}(I)^{\top}\phi_{\text{text}}(P_{k})}{\|\phi_{\text{img}}(I)\|\cdot\|\phi_{\text{text}}(P_{k})\|}.$
The predicted prompt $\hat{T}$ is then selected as the one with the highest similarity score:
$\hat{P}=\arg\max_{P_{k}\in\mathcal{P}}S_{P_{k}},$
and let $S_{\hat{T}}$ denote the scalar similarity score corresponding to the predicted prompt $\hat{T}$.Grad-CAM computes the gradient of the score $S_{c}$ with respect to the feature map $A_{k}(i,j)$, and then applies global average pooling over spatial dimension:
$\alpha_{k}=\frac{1}{H\cdot W}\sum_{i=1}^{H}\sum_{j=1}^{W}\frac{\partial S_{\hat{P}}}{\partial A_{k}(i,j)}.$
The gradient tells how much each feature map channel $A_{k}$ contributes to the model’s prediction $S_{\hat{c}}$. it does this by averaging the gradient of predicted score with respect to all locations $(i,j)$ in that channel and use it act as a weight that reflects the importance of that channel. Based on it, to compute the Grad-CAM heatmap, we first obtain a weighted combination of the activation maps using channel-wise importance weights $\alpha_{k}$, followed by a ReLU non-linearity. Finally, we apply bilinearly upsampling to resize $L$ to the size of original image.
$L_{\text{Grad-CAM}}=\text{Upsample}_{\text{bilinear}}\left(\text{ReLU}\left(\sum_{k=1}^{K}\alpha_{k}\cdot A_{k}\right)\right).$ (1)
@@ 3.3.2 Gradient-Based Attention Rollout for CLIP-ViT
Grad-CAM can be adapted to CLIP-ViT, specifically CLIP [5] with a ViT image encoder. Following the method from [20], instead of using convolutional feature maps, we leverage the self-attention matrices and their gradients to identify input image patches that contribute most to the similarity between an image and a given text prompt. Let $\mathcal{T}^{(l)}\in\mathbb{R}^{N\times N}$ denote the self-attention matrix from the $l$-th transformer block in the ViT encoder, where $N$ is the number of tokens, including the [CLS] token, a special token added at the beginning of the input, and its output representation is used to summarize the entire image for classification tasks. We extract $\mathcal{T}^{(l)}$ after the softmax operation during the forward pass.
@@
$\nabla\mathcal{T}^{(l)}=\frac{\partial S_{\hat{P}}}{\partial\mathcal{T}^{(l)}},$
where $S_{\hat{T}}$ is the predicted similarity score corresponding to the selected text prompt in a zero-shot setting. To capture influence on the similarity score, we compute the gradient-weighted attention map:
@@
Finally, we reshape and upsample the vector $[r_{1},\dots,r_{N-1}]\in\mathbb{R}^{N-1}$ into the input image size using bilinear interpolation. The final heatmap can be constructed as following:
$L_{\text{Grad-CAM-ViT}}=\text{Upsample}_{\text{bilinear}}\left(\text{reshape}\left([r_{1},\dots,r_{N-1}]\right)\right),$ (2)
3.4 Evaluation Metrics
Inspired by [37], we adopt a similar class-aware localization metric in our work to assess vision-language models in surgical settings. In this section, we introduce the metrics for instruments and triplet classification tasks.
3.4.1 Instrument classification metrics
Conventional evaluation for this task in [13] treats it as a multi-label classification problem and reports the F1 score by thresholding the image-text similarity score for each instrument class, selecting the optimal threshold per class, and averaging F1 scores across all instruments.
We now introduce our metrics, let $I$ denote an input image of size $H\times W$, and let $T$ be a corresponding text prompt. The model produces a Grad-CAM heatmap $L(i,j)\in\mathbb{R}^{H\times W}$ using (1) or (2), where each pixel $(i,j)$ reflects the spatial importance of that location for the model’s prediction. We normalize $L$ to the range $[0,1]$, and apply a fixed threshold $\tau\in[0,1]$ to etain information about the "strength" of different regions. We now define the predicted $\tau$-attention region as following:
@@ 3.4.1 Instrument classification metrics
Let $\mathcal{C}=\{c_{1},\dots,c_{K}\}$ denote the set of ground-truth class labels (e.g., surgical instruments) present in the image. In Section 4, we will explain how these classes are used to construct prompts for VLMs using predefined templates. For each class $c_{k}\in\mathcal{C}$, let $\mathcal{G}_{c_{k}}^{(r)}\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ denote the set of pixel coordinates in the $r^{\text{th}}$ annotated region of class $c_{k}$, for $r=1,\dots,N_{k}$. Let $\hat{c}$ be the class predicted by the model. We define the set of all annotated pixels (regardless of class) as:
@@
This ensures that only alignment with correctly predicted classes contributes positively to the score. Beyond conventional classification tasks, understanding the surgical action depicted in an image is also critically important in surgical data. In the instrument classification task, models are not required to predict all classes present in the image. Instead, we generate the heatmap $L_{\text{GradCAM}}$ and $L_{\text{GradCAM-ViT}}$ and evaluate spatial grounding using two metrics: $\tau$-AC, which assesses whether the model attends to any relevant region, and $\tau$-AA, which evaluates whether attention is correctly aligned with the specific object mentioned in the prompt.
3.4.2 Standard Triplet evaluation metrics
(1) Triplet Classification:
To evaluate the performance of vision-language models (VLMs) on surgical action recognition, we adopt conventional classification metrics based on the triplet recognition task introduced in [34]. Each predicted action is represented as a triplet of class labels $\langle\texttt{instrument},\texttt{verb},\texttt{target}\rangle=(s,v,o)$, corresponding to the subject (surgical instrument), verb (action), and object (anatomical target), respectively. Given a ground-truth triplet $(s^{*},v^{*},o^{*})$ and a predicted triplet $(s,v,o)$, we define binary match indicators for each component as follows:
We now formally define the following conventional evaluation metrics we mentioned in section 3.1:
IVT Accuracy (Full Triplet Accuracy): $\text{IVT}=\mathbbm{1}_{(s,v,o)=(s^{*},v^{*},o^{*})}$
IV Accuracy (Instrument-Verb Accuracy): $\text{IV}=\mathbbm{1}_{(s,v)=(s^{*},v^{*})}$
IT Accuracy (Instrument-Target Accuracy): $\text{IT}=\mathbbm{1}_{(s,o)=(s^{*},o^{*})}$
These combinations are evaluated because surgical actions are inherently instrument-centric; both the verb and target gain semantic meaning only when grounded in the context of the instrument performing the action.
@@ IV Accuracy (Instrument-Verb Accuracy): $\text{IV}=\mathbbm{1}_{(s,v)=(s^{*},v^{*})}$
In the triplet classification task, the traditional and straightforward way of evaluating performance is to compute the F1 classification score for each class. However, in this paper, beyond assessing classification accuracy, we also challenge the spatial reasoning ability of VLMs. However, when it comes to localizing the bounding box of an action verb, the problem becomes more complex. As discussed in [25], the localization of the action region is closely related to the instrument performing the action. In particular, when an instrument is actively performing an action, the action is spatially grounded near the location of the instrument tip. Due to the lack of labeled action verb bounding boxes, we propose the Action Reasoning Score (ARS), it measures whether the model’s attention for a predicted action (verb) overlaps with the bounding box of the predicted instrument (subject). Let a predicted triplet be denoted by $(s,v,o)$, where $s\in\mathcal{S}$ is the subject, $v\in\mathcal{V}$ is the verb (e.g., an action such as cut), and $o\in\mathcal{O}$ is the object (e.g., tissue) and ground-truth triplet is $(s^{*},v^{*},o^{*}).$We compute ARS only when: $s=s^{*}\quad\text{and}\quad v=v^{*}.$ Let the image be defined over a pixel grid of height $H$ and width $W$, with each pixel indexed by coordinates $(i,j)\in\{1,\dots,H\}\times\{1,\dots,W\}$. Let $L_{v}(i,j)\in[0,1]$ denote the value of Grad-CAM attention produced by (1) or (2) at pixel $(i,j)$ for the predicted verb $v$. We will later show how to generate the attention heatmap $L_{v}$ and construct prompts using predefined templates for VLMs in section 4.
@@
Note that this score provides a soft measure of how much the verb-specific attention overlaps with the instrument’s spatial region, offering insight into whether the model is focusing on a plausible area for grounding the action. While there is no strict threshold for overlapping, a higher value generally indicates better spatial alignment, whereas a near-zero value suggests the model is likely attending to an irrelevant region.
@@ 4 Experiments Set up
All experiments use a fixed Grad-CAM threshold of $\tau=0.3$ to generate attention maps for grounding evaluation. We first introduce the prompt designs used across all tasks, followed by a description of the benchmarked models evaluated in this work. We evaluate our methods using three surgical video datasets. The Cholec80-BBox dataset [26, 27], which extends the original Cholec80 dataset [10], consists of five laparoscopic cholecystectomy videos with instrument bounding box annotations and is used for instrument classification. The CoPESD dataset [11] includes two endoscopic submucosal dissection procedures with two type of tool annotations and is also used for instrument classification. The CholecT45 dataset [28] comprises 45 laparoscopic cholecystectomy videos annotated with action triplets and instrument bounding boxes, and is used for evaluating both triplet classification and action reasoning.
@@
To align with the prompt-based input format of VLMs, we construct a pool of prompts using template-based formulations. For instrument classification, we use prompts template: “an image showing a {instrument class} in use”, covering all possible instrument classes in Cholec80. For triplet-based action reasoning, we use the template “I use a {instrument} to {verb} the {target}.
For triplet reasoning, we compute the Action Reasoning Score (ARS) is computed only for cases where $\text{IV}=1$, that is, both the subject (instrument) and verb predictions are correct. In our implementation, similar to the instrument classification setup, we first generate a pool of triplet prompts using the template “I use a {instrument} to {verb} the {target}, covering all possible instrument, verb, and target classes. We then identify frames where both the predicted instrument and verb match the ground truth. For these correctly classified frames, we reuse the predicted verb to construct a secondary prompt of the form “I am performing {predicted_verb}. Finally, we generate its Grad-CAM heatmap $L_{v}$ and compute the action reasoning scores (6).
@@ 4.1 Benchmarked Models
In this work, we evaluate four VLMs: CLIP-ViT[5], HecVL[8], PeskaVLP[7], and SurgVLP[9]. CLIP-ViT [5] is the classic and widely adopted architecture for multi-modal learning, trained on large-scale natural image-text pairs. CLIP has demonstrated strong zero-shot performance across various domains, making it a natural reference point for evaluating domain transfer to surgical imagery.
The other three models: SurgVLP, PeskaVLP, and HecVL are trained using surgical video datasets with varying levels of supervision, and architectural adaptations. These models are designed specifically to address the challenges of understanding visual-linguistic information in surgical environments, such as tool usage, procedural context, and anatomy. We aim to assess not only a deeper understanding of the reasoning capabilities of VLMs in surgical classification tasks, but also the limitations and trade-offs between general-purpose and task-specific VLMs within the current surgical domain.
Figure 2: Examples of triplet classification tasks using CLIP-ViT with Grad-CAM visualize the verb heatmap $L_{v}$, where CLIP-ViT successfully classifies all images in Panels (a) and (b) with the correct instrument and verb. Green bounding boxes indicate ground truth annotation from video 42 and 43 in CholecT80-BBox dataset. Panels (a) illustrate Ideal action reasoning, with attention focused near the instrument and high ARS scores. In contrast, panels (b) show failure cases: despite correct triplet predictions, the attention is misplaced on background tissues, leading to ARS scores of 0.00. These examples highlight that some correct classifications are not based on the correct spatial evidences.
@@ 5.1 Instrument Classification Performance
Table 1 shows the spatial reasoning performance of five vision-language models across two datasets. Overall, all VLMs tested so far struggle under complex surgical environments. CLIP achieves the highest scores on the CoPESD dataset, which contains fewer visible instrument instances. It consistently outperforms the other models in both $\tau$-AC and $\tau$-AA metrics across most videos, achieving a $\tau$-AC of 0.3405 and $\tau$-AA of 0.0680 on the CoPESD dataset, and a $\tau$-AC of 0.2395 and $\tau$-AA of 0.1021 on the Cholec8k Bounding Box dataset. In contrast, despite being trained on surgical lectures, HecVL, PeskaVLP, and SurgVLP perform worse compared to the general-purpose model CLIP. We hypothesize that this is because these models are primarily designed to understand high-level surgical phases rather than to perform fine-grained, frame-level instrument classification. Overall, while VLMs demonstrate some ability to localize surgical instruments, they still lack the capacity to consistently and correctly classify them in challenging surgical scenes.